A mass of 9kg draws a second mass of 4.5 kg up a smooth 30° incline by a light cord passing over a smooth pulley. Find the velocities of the two bodies after they have moved a distance of 3m from rest(Use conservation of energy)


Clearly show the following values:


Direction of movement=??

Friction force=??

Ek=??

Ep=??

ETR=??

ETE=??

EFRICTION=??

Energy Equation=??

VA=??

VB=??



1
Expert's answer
2021-09-29T04:37:36-0400

Direction of movement:

Mass A moves upwards, mass B moves downwards

Frictional force; this is assumed to be zero since the surface is smooth

Ek=

Ep=

Energy conservation requires that the change in potential energy of the system m1+m2 is of same magnitude to change in kinetic energy of the system but of a negative sign,

∆Kin=-∆Pout

The two masses are connected and move with the same speed V,

Efriction=0

Energy equation;

"\frac{1}{2}m_1V^{2}+\frac{1}{2}m_2V^{2}=(m_1g∆h_1+m_2g∆h_2+0)"

"(m_1+m_2)V^{2}=2(m_1gh_2Sin\theta+m_2gh_2)"

"V^{2}=\frac{2(m_1gh_2Sin\theta+m_2gh_2)}{(m_1+m_2)}"

"V=\sqrt\frac{2(m_1gh_2Sin\theta+m_2gh_2)}{(m_1+m_2)}"

"V=\sqrt\frac{2(4.5*9.81*3Sin30°+9*9.81*3)}{(4.5+9)}"

"V=7.0036m\/s"

"V_A=7.0036m\/s"

"V_B=7.0036m\/s"